OpenStudy (zoiedaniella):

Anyone mind helping me understand this? Given the functions a(x) = 4x2 + 2x − 3 and b(x) = x − 1, identify the oblique asymptote of the function a of x over the function b of x

OpenStudy (michele_laino):

we have to write the oblique asymptotes of this function: \[f\left( x \right) = \frac{{a\left( x \right)}}{{b\left( x \right)}} = \frac{{4{x^2} + 2x - 3}}{{x - 1}}\]

OpenStudy (welshfella):

Do the long division - what do you get?

OpenStudy (michele_laino):

first step we have to evaluate this limit: \[\mathop {\lim }\limits_{x \to \pm \infty } \frac{{f\left( x \right)}}{x} = \mathop {\lim }\limits_{x \to \pm \infty } \frac{{4{x^2} + 2x - 3}}{{x\left( {x - 1} \right)}} = ...?\]

OpenStudy (zoiedaniella):

that will be 4x+6+3/x-1 simplified correct?

OpenStudy (michele_laino):

hint: divide numerator an denominator by \(x\)

OpenStudy (zoiedaniella):

mmkay, lemme do this on paper real quick, brb.

OpenStudy (michele_laino):

starting from your result, of course

OpenStudy (zoiedaniella):

okay, do you mind writing out what I would divide? I just really confused myself for a min there.

OpenStudy (michele_laino):

mean this: starting from your expression: (4x+2-(3/x))/(1-(1/x)) please divide again numerator and denominator by x

OpenStudy (michele_laino):

sorry, please start from this function: \[\frac{{4{x^2} + 2x - 3}}{{{x^2} - x}}\] and divide numerator and enominator by \(x^2\)

OpenStudy (michele_laino):

denominator*

OpenStudy (zoiedaniella):

that would just cancel out the x^2's correct?

OpenStudy (zoiedaniella):

so it would just be 4+2x-3/-x ?

OpenStudy (zoiedaniella):

and now do long division?

OpenStudy (michele_laino):

I got this: for numerator \[4{x^2} + 2x - 3 \to 4 + \frac{2}{x} - \frac{3}{{{x^2}}}\]

OpenStudy (michele_laino):

now, do the same with the denominator please

OpenStudy (zoiedaniella):

\[x^2/x^2 -x/x^2 \] ?? i'm sorry i'm very confused haha

OpenStudy (michele_laino):

correct! SO, after a simplification we get: \[\frac{{{x^2} - x}}{{{x^2}}} = 1 - \frac{1}{x}\]

OpenStudy (michele_laino):

oops.. So*...

OpenStudy (zoiedaniella):

ohhhh okay

OpenStudy (michele_laino):

now we have to evaluate this limit: \[\mathop {\lim }\limits_{x \to \pm \infty } \frac{{f\left( x \right)}}{x} = \mathop {\lim }\limits_{x \to \pm \infty } \frac{{4{x^2} + 2x - 3}}{{x\left( {x - 1} \right)}} = \mathop {\lim }\limits_{x \to \pm \infty } \frac{{4 + \frac{2}{x} - \frac{3}{{{x^2}}}}}{{1 - \frac{1}{x}}} = ...?\]

OpenStudy (michele_laino):

hint: we have this: as \(x\) goes to infinity, the quantity \(1/x\) goes to zero

OpenStudy (michele_laino):

and also the quantity \(1/x^2\) goes to zero

OpenStudy (zoiedaniella):

doing it on paper, one min. So the answer would be \[4+(6x-3)/x^2-x\]

OpenStudy (michele_laino):

the answer is: